The ball rolls off the stairway with a horizontal velocity of magnitude 5m/s . The steps are 8cm high and 8cm wide which step will the ball hit first

Grade:11

1 Answers

Arun
6 years ago
Horizontal distance covered by ball =nb
Vertical distance covered by ball =nh
If the time taken to hit the edge of nth step is t, then
nb = ut
nh = 1/2 gt^2
Eliminating t, we get
n = 2h* u^2/ g*b^2
Now in the question we have to find the value of n when,
h = 8, u= 5, b =8
So
n = 2*8*25*100/64*10 = 4000/64 = 62.5
Hence it will hit 62th steps